Assume that banks lend out all their excess reserves. Currently, the legal reserves that banks must hold equal $11.5 billion. If the Federal Reserve decreases its reserve requirement from 10% to 5%, then there is potential for the whole banking system to raise money supply by:

a) $11.5 billion

b) $230 billion

c) $115 billion

d) $57.5 billion

e) $575 billion
Answer #1

c) $115 billion

(As the reserve requirement is halved, so new reserves are 11.5/2 = 5.75 billion
So, change in deposits = 5.75 billion
Change in money supply = (1/reserve ratio)*change in deposits = (1/5%)*5.75 = (1/0.05)*5.75 = (20)*(5.75) = 115 billion)